# Científicas (CSIC) and Universitat de València (UV), 46980 Valencia, Spain

### v1

@@ 3.1.1 Higgs self-energy and self-coupling
The left panel of Fig. 5 shows the on- and off-pole 68% CL projected sensitivities at FCC-ee in the $C_{H\Box}$ vs. $C_{H}$ plane, that modify the Higgs 2-point and 3-point functions respectively. The blue-shaded ellipse corresponds to the off-pole constraints (at LO for $C_{H\Box}$ and NLO for $C_{H}$) while the orange-shaded region denotes the on-pole bounds (at NLO for $C_{H\Box}$ and NNLO for $C_{H}$). We note that an otherwise flat direction for the off-pole observables is broken by measuring $\sigma(e^{+}e^{-}\to ZH)$ at 240 and 365 GeV. The on-pole observables then further constrain an almost orthogonal direction, leading to a clear complementarity in the combination shown by the green ellipse. While in general the region probed by on-pole observables depends on the inclusion of the quadratic terms in Eq. 12, the combination with off-pole $ZH$ data clearly limits $C_{H}$ to the linear regime represented by the region between dashed orange lines.
@@ 3.1.2 Higgs-vector-vector couplings
We now turn to the right panel of Fig. 5 which shows the on- and off-pole 68% CL projected sensitivities at FCC-ee in the $C_{HW}$ vs. $C_{HB}$ plane that modify the 3-point $hVV$ vertex. The complementarity is again evident between the on-pole constraints in the orange-shaded region and the off-pole blue-shaded regions. Both on- and off-pole observables have flat directions that are lifted by the combination to give the combined green elliptical region.
@@ 3.1.3 Higgs sector combined fit
We report here a combined on- and off-pole likelihood for $C_{H},C_{H\Box},C_{HW},C_{HB}$ from a 4-parameter Gaussian fit. The 1-$\sigma$ intervals for the Wilson coefficients renormalised at 1 TeV are
$\left(\begin{array}[]{cc}C_{H}\\ C_{H\Box}\\ C_{HW}\\ C_{HB}\\ \end{array}\right)=\pm\left(\begin{array}[]{cc}1.089\\ 0.052\\ 0.020\\ 0.034\\ \end{array}\right)\,,$ (13)
@@
$\left(\begin{array}[]{cccc}1&0.107&0.763&-0.637\\ 0.107&1&-0.137&-0.697\\ 0.763&-0.137&1&-0.592\\ -0.637&-0.697&-0.592&1\\ \end{array}\right)\,.$ (14)
We note that the combination of on-pole and off-pole data here enables a sensible closed fit to be obtained for this 4-parameter subset of Higgs coupling operators.
@@ 3.2 Pure gauge operators
Of the 4 oblique parameters $\hat{S}$, $\hat{T}$, $\hat{W}$, and $\hat{Y}$ that correspond to dimension-6 operators, it is thought that only $\hat{S}$ and $\hat{T}$ are strongly constrained on the $Z$ pole while $\hat{W}$ and $\hat{Y}$ are best constrained at higher energies due to their energy enhancement. However, as argued in the introduction, unlike at LEP this is no longer expected to be the case at FCC-ee, where the statistical improvement in accuracy at the $Z$ pole can compensate for the energy enhancement of off-pole processes despite both being at leading order. For example, in $WW$ processes this is due to the relative improvement in the number of $WW$ and $Z$ boson pairs in going from LEP to FCC-ee, $\left(N_{Z}/N_{WW}\right)_{FCC}\sim 10^{5}\gg\left(N_{Z}/N_{WW}\right)_{LEP}\sim 10^{2}$.
@@ 3.2.1 Oblique parameters $\hat{S}$, $\hat{T}$, $\hat{W}$, $\hat{Y}$
As is well known, the $\hat{S}$ and $\hat{T}$ parameters are strongly constrained by the pole observables. On the other hand, $\hat{W}$ and $\hat{Y}$ receive competitive constraints both on and off the pole, as shown recently Greljo:2024ytg . This is expected from our earlier accuracy complements energy discussion as being due to a competition between the energy enhancement these parameters receive off pole versus the high statistics on pole. We have checked that it is an excellent approximation to neglect the running of these 4 operators between the off-pole energies down to $m_{Z}$. We therefore treat the Wilson coefficients entering the map to $\hat{S},\hat{T},\hat{W},\hat{Y}$ as evaluated at $m_{Z}$ and do the combined on-and-off pole fit. We find that all the oblique parameters can be constrained at the $10^{-5}$ level
$\left(\begin{array}[]{cc}\hat{S}\\ \hat{T}\\ \hat{W}\\ \hat{Y}\\ \end{array}\right)=\pm\left(\begin{array}[]{cc}2.907\\ 1.264\\ 0.681\\ 2.583\\ \end{array}\right)\times 10^{-5}\,,$ (22)
@@
$\rho=\left(\begin{array}[]{cccc}1.&0.811&0.336&0.919\\ 0.811&1.&0.096&0.578\\ 0.336&0.096&1.&0.238\\ 0.919&0.578&0.238&1.\\ \end{array}\right)\,.$ (23)
These results are in good agreement with Ref. Greljo:2024ytg . Having done the oblique parameter fit for completeness, in what follows we will drop the operators corresponding to $\hat{S}$ and $\hat{T}$ as any new physics model generating sizeable contributions to them will be dominated by on-pole phenomenology. In that case, one can simply use the likelihood given by Eqs. 22 and 23. However in general it is important to note that the oblique parameters only correctly capture on-pole physics for flavour-universal theories, and only at leading order. For non-universal theories or when the pole observables are computed at NLO, there are NP contributions that cannot be absorbed into the oblique parameters, and the SMEFT framework is required.
@@ 3.2.3 Gauge sector combined fit
$\left(\begin{array}[]{cc}S_{2W}\\ S_{2B}\\ S_{3W}\\ S_{HW}\\ S_{HB}\\ \end{array}\right)=\pm\left(\begin{array}[]{cc}0.10\\ 0.29\\ 1.99\\ 2.78\\ 2.47\\ \end{array}\right)\times 10^{-2}\,,$ (29)
@@
$\rho=\left(\begin{array}[]{ccccc}1.&0.08&-0.36&-0.05&-0.10\\ 0.08&1.&-0.78&-0.47&0.26\\ -0.36&-0.78&1.&-0.01&0.15\\ -0.05&-0.47&-0.01&1.&-0.70\\ -0.10&0.26&0.15&-0.70&1.\\ \end{array}\right)\,.$ (30)
We plot in Fig. 8 the on-pole and off-pole projected 68% CL sensitivities at FCC-ee in the plane of $S_{2W}$ vs. $S_{2B}$, corresponding to the $W$ and $Y$ parameters, on the left and for $S_{2W}$ vs $S_{3W}$ on the right. We see overall comparable constraints in the 2-dimensional ellipses, with off(on)-pole data giving stronger bounds along the $S_{2W}$ ($S_{2B}$) direction and on-pole data being crucial for restricting the overall combination to the green ellipse.
@@
The on- and off-pole sensitivity to four-fermion (4F) operators has already received significant attention in the literature. For example, the $Z$-pole sensitivity to 4F operators, especially those involving top quarks, has been studied in Dawson:2019clf ; Dawson:2022bxd ; Allwicher:2023aql ; Garosi:2023yxg ; Allwicher:2023shc ; Bellafronte:2023amz ; Stefanek:2024kds ; Haisch:2024wnw ; Ge:2024pfn . On the other hand, off-pole sensitivity in the $e^{+}e^{-}\rightarrow\bar{f}f$ processes at the $WW$, $ZH$, and $t\bar{t}$ energies was studied recently in Ref. Greljo:2024ytg via a dedicated flavor tagging analysis. Here we review and summarise the main conclusions. The projected sensitivities are shown in Fig. 1 for the four-quark (4Q), four-lepton (4L), and two-quark-two-lepton (2Q2L) operators, with darker shading indicating a stronger sensitivity to the effective scale of new physics $\Lambda^{i}_{\rm eff}=(\mathcal{C}_{i})^{-1/2}$, while Fig. 9 summarizes the relevant diagrams.
$e^{+}$$e^{-}$
@@
For the semi-leptonic (2Q2L) operators, we see better sensitivity from on-pole data for operators involving top quarks and no electrons, namely: $[\mathcal{C}^{(1)}_{lq}]_{3333}$, $[\mathcal{C}^{(3)}_{lq}]_{3333}$, $[\mathcal{C}_{lu}]_{3333}$, $[\mathcal{C}_{eu}]_{3333}$, and $[\mathcal{C}_{qe}]_{3333}$. These operators run strongly into EW vertex corrections $\propto N_{c}y_{t}^{2}$ by closing the top loop and attaching two Higgses. On the other hand, off-pole data unsurprisingly performs better for operators involving electrons as they enter at LO and receive an energy enhancement. Still, a competitive bound can be achieved at the $Z$-pole if the operator has top quarks such as: $[\mathcal{C}^{(1)}_{lq}]_{1133}$, $[\mathcal{C}^{(3)}_{lq}]_{1133}$, $[\mathcal{C}_{lu}]_{1133}$, $[\mathcal{C}_{eu}]_{1133}$, and $[\mathcal{C}_{qe}]_{3311}$. Finally, other semi-leptonic operators without tops or electrons contribute to both on- and off-pole observables only at the loop level (see Fig. 9), typically with comparable or slightly better sensitivity on the pole.
@@
Four quark operators enter via the loop diagrams in Fig. 9 for both on-pole and off-pole measurements, with on-pole data providing better sensitivity for operators involving top quarks such as $[\mathcal{C}^{(1)}_{qq}]_{3333}$, $[\mathcal{C}^{(1)}_{qu}]_{3333}$, and $[\mathcal{C}_{uu}]_{3333}$. [^3] In the case of $[\mathcal{C}_{uu}]_{3333}$, the contribution to $Z$-pole observables occurs starting at the 2-loop level, but it is nonetheless phenomenologically relevant Allwicher:2023aql ; Allwicher:2023shc ; Stefanek:2024kds . In the absence of tops, the dominant contribution comes from EW gauge running both on and off the pole. In this case, the constraints are typically comparable which can be understood by considering the energy enhancement off-pole vs. statistics on-pole, again demonstrating the principle that accuracy complements energy.
@@
Finally, four-lepton operators contribute at leading order to $e^{+}e^{-}\rightarrow\bar{f}f$ if they involve electrons, in which case these observables dominate. A partial exception is $[\mathcal{C}_{ll}]_{1122}$ which also receives a noteable on-pole constraint due to its running into $[\mathcal{C}_{ll}]_{1221}$. Otherwise, we are again in the situation of loop contributions via gauge running, leading to comparable constraints. In particular, this includes the pure third-generation case with operators such as $[\mathcal{C}_{ll}]_{3333}$, $[\mathcal{C}_{ee}]_{3333}$ and $[\mathcal{C}_{le}]_{3333}$.
@@ 4 Specific UV model examples
Having illustrated the principle that accuracy complements energy in the SMEFT, we now move to examples in concrete UV-complete models. While we focus here on bosonic models, example models showing the power of combining on- and off-pole sensitivity to four fermion operators can be found in Ref. Greljo:2024ytg .
@@ 4.1 Real singlet scalar
Figure 10: FCC-ee sensitivity at 68% CL on the real singlet scalar model. The dashed lines show the limits if only $\mathcal{C}_{H\Box}$ is matched, which neglects the correlated contribution from $\mathcal{C}_{H}$. For parameters in the red region, the model can give rise to a first-order EW phase transition (curve taken from Crawford:2024nun ).
In Fig. 10, we show the on- and off-pole constraints on the real singlet scalar model expected at FCC-ee, both of which exclude a first-order electroweak phase transition for the mass range shown. In particular, virtually all the loryon parameter space can now be probed by the inclusion of on-pole data that significantly extends the off-pole bounds. The dashed blue line shows the off-pole constraints assuming only $\mathcal{C}_{H\Box}$ in the fit (for comparison to previous analyses), while the solid blue line is the full result also including $\mathcal{C}_{H}$. The weaker constraint from the full result can be understood as follows. First, we have $\mathcal{C}_{H}/\mathcal{C}_{H\Box}=2\kappa$, so the importance of $\mathcal{C}_{H}$ relative to $\mathcal{C}_{H\Box}$ grows linearly with $\kappa$. Next, these two Wilson coefficients contribute to the $ZH$ cross section with the ratio $\mathcal{C}_{H}/\mathcal{C}_{H\Box}\approx-0.07$. This means that for $\kappa\approx 7$, the NLO contribution of $\mathcal{C}_{H}$ becomes of the same size as the LO contribution of $\mathcal{C}_{H\Box}$ and there is a partial cancellation that leads to the weakened bound shown in Fig. 10.

### v2

@@ 3.1.1 Higgs self-energy and self-coupling
The left panel of Fig. 5 shows the on- and above-pole 68% CL projected sensitivities at FCC-ee in the $C_{H\Box}$ vs. $C_{H}$ plane, that modify the Higgs 2-point and 3-point functions respectively. The blue-shaded ellipse corresponds to the above-pole constraints (at LO for $C_{H\Box}$ and NLO for $C_{H}$) while the orange-shaded region denotes the on-pole bounds (at NLO for $C_{H\Box}$ and NNLO for $C_{H}$). We note that an otherwise flat direction for the above-pole observables is broken by measuring $\sigma(e^{+}e^{-}\to ZH)$ at 240 and 365 GeV. The on-pole observables then further constrain an almost orthogonal direction, leading to a clear complementarity in the combination shown by the green ellipse. While in general the region probed by on-pole observables depends on the inclusion of the quadratic terms in Eq. 13, the combination with above-pole $ZH$ data clearly limits $C_{H}$ to the linear regime represented by the region between dashed orange lines.
@@ 3.1.2 Higgs-vector-vector couplings
We now turn to the right panel of Fig. 5 which shows the on- and above-pole 68% CL projected sensitivities at FCC-ee in the $C_{HW}$ vs. $C_{HB}$ plane that modify the 3-point $hVV$ vertex. The complementarity is again evident between the on-pole constraints in the orange-shaded region and the above-pole blue-shaded regions. Both on- and above-pole observables have flat directions that are lifted by the combination to give the combined green elliptical region.
@@ 3.1.3 Higgs sector combined fit
We report here a combined on- and above-pole likelihood for $C_{H},C_{H\Box},C_{HW},C_{HB}$ from a 4-parameter Gaussian fit. The 1-$\sigma$ intervals for the Wilson coefficients renormalised at 1 TeV are
$\left(\begin{array}[]{cc}C_{H}\\ C_{H\Box}\\ C_{HW}\\ C_{HB}\\ \end{array}\right)=\pm\left(\begin{array}[]{cc}0.932\\ 0.052\\ 0.015\\ 0.031\\ \end{array}\right)\,$, (14)
@@
$\left(\begin{array}[]{cccc}1&0.174&0.772&-0.612\\ 0.174&1&-0.122&-0.786\\ 0.772&-0.122&1&-0.489\\ -0.612&-0.786&-0.489&1\\ \end{array}\right)\,.$ (15)
We note that the combination of on-pole and above-pole data here enables a sensible closed fit to be obtained for this 4-parameter subset of Higgs coupling operators.
@@ 3.2 Pure gauge operators
Of the 4 oblique parameters $\hat{S}$, $\hat{T}$, $\hat{W}$, and $\hat{Y}$ that correspond to dimension-6 operators, it is thought that only $\hat{S}$ and $\hat{T}$ are strongly constrained on the $Z$ pole while $\hat{W}$ and $\hat{Y}$ are best constrained at higher energies due to their energy enhancement. However, as argued in the introduction, unlike at LEP this is no longer expected to be the case at FCC-ee, where the statistical improvement in accuracy at the $Z$ pole can compensate for the energy enhancement of above-pole processes despite both being at leading order. For example, in $WW$ processes this is due to the relative improvement in the number of $WW$ and $Z$ boson pairs in going from LEP to FCC-ee, $\left(N_{Z}/N_{WW}\right)_{FCC}\sim 10^{5}\gg\left(N_{Z}/N_{WW}\right)_{LEP}\sim 10^{2}$.
@@ 3.2.1 Oblique parameters $\hat{S}$, $\hat{T}$, $\hat{W}$, $\hat{Y}$
As is well known, the $\hat{S}$ and $\hat{T}$ parameters are strongly constrained by the pole observables. On the other hand, $\hat{W}$ and $\hat{Y}$ receive competitive constraints both on and off the pole, as shown recently Greljo:2024ytg . This is expected from our earlier accuracy complements energy discussion as being due to a competition between the energy enhancement these parameters receive above-pole versus the high statistics on pole. We have checked that it is an excellent approximation to neglect the running of these 4 operators between the above-pole energies down to $m_{Z}$. We therefore treat the Wilson coefficients entering the map to $\hat{S},\hat{T},\hat{W},\hat{Y}$ as evaluated at $m_{Z}$ and do the combined on-and-above pole fit. We find that all the oblique parameters can be constrained at the $10^{-5}$ level
$\left(\begin{array}[]{cc}\hat{S}\\ \hat{T}\\ \hat{W}\\ \hat{Y}\\ \end{array}\right)=\pm\left(\begin{array}[]{cc}2.26\\ 1.16\\ 0.49\\ 1.90\\ \end{array}\right)\times 10^{-5}\,$, (23)
@@
$\rho=\left(\begin{array}[]{cccc}1&0.795&0.331&0.868\\ 0.795&1&0.088&0.461\\ 0.331&0.088&1&0.256\\ 0.868&0.461&0.256&1\\ \end{array}\right)\,.$ (24)
These results are in good agreement with Ref. Greljo:2024ytg . Having done the oblique parameter fit for completeness, in what follows we will drop the operators corresponding to $\hat{S}$ and $\hat{T}$ as any new physics model generating sizeable contributions to them will be dominated by on-pole phenomenology. In that case, one can simply use the likelihood given by Eqs. 23 and 24. However in general it is important to note that the oblique parameters only correctly capture on-pole physics for flavour-universal theories, and only at leading order. For non-universal theories or when the pole observables are computed at NLO, there are NP contributions that cannot be absorbed into the oblique parameters, and the SMEFT framework is required.
@@ 3.2.3 Gauge sector combined fit
$\left(\begin{array}[]{cc}S_{2W}\\ S_{2B}\\ S_{3W}\\ S_{HW}\\ S_{HB}\\ \end{array}\right)=\pm\left(\begin{array}[]{cc}0.08\\ 0.22\\ 1.52\\ 2.23\\ 2.34\\ \end{array}\right)\times 10^{-2}\,$, (30)
@@
$\rho=\left(\begin{array}[]{ccccc}1&0.10&-0.34&-0.07&-0.06\\ 0.10&1&-0.74&-0.42&0.16\\ -0.34&-0.74&1&-0.09&0.28\\ -0.07&-0.42&-0.09&1&-0.69\\ -0.06&0.16&0.28&-0.69&1\\ \end{array}\right)\,.$ (31)
We plot in Fig. 8 the on-pole and above-pole projected 68% CL sensitivities at FCC-ee in the plane of $S_{2W}$ vs. $S_{2B}$, corresponding to the $W$ and $Y$ parameters, on the left and for $S_{2W}$ vs $S_{3W}$ on the right. We see overall comparable constraints in the 2-dimensional ellipses, with off(on)-pole data giving stronger bounds along the $S_{2W}$ ($S_{2B}$) direction and on-pole data being crucial for restricting the overall combination to the green ellipse.
@@
The on- and above-pole sensitivity to four-fermion (4F) operators has already received significant attention in the literature. For example, the $Z$-pole sensitivity to 4F operators, especially those involving top quarks, has been studied in Dawson:2019clf ; Dawson:2022bxd ; Allwicher:2023aql ; Garosi:2023yxg ; Allwicher:2023shc ; Bellafronte:2023amz ; Stefanek:2024kds ; Haisch:2024wnw ; Ge:2024pfn . On the other hand, above-pole sensitivity in the $e^{+}e^{-}\rightarrow\bar{f}f$ processes at the $WW$, ZHZH, and $t\bar{t}$ energies was studied recently in Ref. Greljo:2024ytg via a dedicated flavor tagging analysis. Here we review and summarise the main conclusions. The projected sensitivities are shown in Fig. 1 for the four-quark (4Q), four-lepton (4L), and two-quark-two-lepton (2Q2L) operators, with darker shading indicating a stronger sensitivity to the effective scale of new physics $\Lambda^{i}_{\rm eff}=(\mathcal{C}_{i})^{-1/2}$, while Fig. 9 summarizes the relevant diagrams.
$^{+}ee^{-}$
@@
For the semi-leptonic (2Q2L) operators, we see better sensitivity from on-pole data for operators involving top quarks and no electrons, namely: $\mathcal{C}^{(1)}_{lq}]_{3333}$, $\mathcal{C}^{(3)}_{lq}]_{3333}$, $\mathcal{C}_{lu}]_{3333}$, $\mathcal{C}_{eu}]_{3333}$, and $\mathcal{C}_{qe}]_{3333}$. These operators run strongly into EW vertex corrections $\propto N_{c}y_{t}^{2}$ by closing the top loop and attaching two Higgses. On the other hand, above-pole data unsurprisingly performs better for operators involving electrons as they enter at LO and receive an energy enhancement. Still, a competitive bound can be achieved at the $Z$-pole if the operator has top quarks such as: $\mathcal{C}^{(1)}_{lq}]_{1133}$, $\mathcal{C}^{(3)}_{lq}]_{1133}$, $\mathcal{C}_{lu}]_{1133}$, $\mathcal{C}_{eu}]_{1133}$, and $\mathcal{C}_{qe}]_{3311}$. Finally, other semi-leptonic operators without tops or electrons contribute to both on- and above-pole observables only at the loop level (see Fig. 9), typically with comparable or slightly better sensitivity on the pole.
@@
Four quark operators enter via the loop diagrams in Fig. 9 for both on-pole and above-pole measurements, with on-pole data providing better sensitivity for operators involving top quarks such as $\mathcal{C}^{(1)}_{qq}]_{3333}$, $\mathcal{C}^{(1)}_{qu}]_{3333}$, and $\mathcal{C}_{uu}]_{3333}$. [^3] In the case of $\mathcal{C}_{uu}]_{3333}$, the contribution to $Z$-pole observables occurs starting at the 2-loop level, but it is nonetheless phenomenologically relevant Allwicher:2023aql ; Allwicher:2023shc ; Stefanek:2024kds . In the absence of tops, the dominant contribution comes from EW gauge running both on and off the pole. In this case, the constraints are typically comparable which can be understood by considering the energy enhancement above-pole vs. statistics on-pole, again demonstrating the principle that accuracy complements energy.
@@
Finally, four-lepton operators contribute at leading order to $e^{+}e^{-}\rightarrow\bar{f}f$ if they involve electrons, in which case these observables dominate. A partial exception is $\mathcal{C}_{ll}]_{1122}$ which also receives a noteable on-pole constraint due to its running into $\mathcal{C}_{ll}]_{1221}$. Otherwise, we are again in the situation of loop contributions via gauge running, leading to comparable constraints. In particular, this includes the pure third-generation case with operators such as $\mathcal{C}_{ll}]_{3333}$, $\mathcal{C}_{ee}]_{3333}$ and $\mathcal{C}_{le}]_{3333}$.
@@ 4 Specific UV model examples
Having illustrated the principle that accuracy complements energy in the SMEFT, we now move to examples in concrete UV-complete models. While we focus here on bosonic models, example models showing the power of combining on- and above-pole sensitivity to four fermion operators can be found in Ref. Greljo:2024ytg .
@@ 4.1 Real singlet scalar
Figure 10: FCC-$e$ sensitivity at 68% CL on the real singlet scalar model. The dashed lines show the limits if only $\mathcal{C}_{H\Box}$ is matched, which neglects the correlated contribution from $\mathcal{C}_{H}$. For parameters in the red region, the model can give rise to a first-order EW phase transition (curve taken from Crawford:2024nun ).
In Fig. 10, we show the on- and above-pole constraints on the real singlet scalar model expected at FCC-ee, both of which exclude a first-order electroweak phase transition for the mass range shown. In Fig. 10, we show the constraints on the real singlet scalar model expected at FCC-ee, running from a fixed renormalization scale $\Lambda=$500\text{\,}\mathrm{GeV}$\$ for simplicity. Both the on- and above-pole sets of observables would exclude a first-order electroweak phase transition for the mass range displayed. In particular, virtually all the loryon parameter space can now be probed by the inclusion of on-pole data that significantly extends the above-pole bounds. The dashed blue line shows the above-pole constraints assuming only $\mathcal{C}_{H\Box}$ in the fit (for comparison to previous analyses), while the solid blue line is the full result also including $\mathcal{C}_{H}$. The weaker constraint from the full result can be understood as follows. First, we have $\mathcal{C}_{H}/\mathcal{C}_{H\Box}=2\kappa$, so the importance of $\mathcal{C}_{H}$ relative to $\mathcal{C}_{H\Box}$ grows linearly with $\kappa$. Next, these two Wilson coefficients contribute to the $ZH$ cross section with the ratio $\mathcal{C}_{H}/\mathcal{C}_{H\Box}\approx-0.07$. This means that for $\kappa\approx 7$, the NLO contribution of $\mathcal{C}_{H}$ becomes of the same size as the LO contribution of $\mathcal{C}_{H\Box}$ and there is a partial cancellation that leads to the weakened bound shown in Fig. 10.